n insurance policy reimburses 100% for losses up to $100, less a deductible. In addition, the policy reimburses 50% of losses beyond $100. The deductible is $20 and losses follow an Exponential distribution with mean $80. Calculate the probability that the reimbursement for a loss is less than $100, given that (12) the reimbursement is greater than SO A) 0.202 B) 0.632 (C)0.736 D) 0.777 E) 0.845 The circled answer is wrong please show the work to arrive at a correct answer please.
Answer #1

here probability reimbursement is greater then 0 =P(X>20)=e-x/\beta =e-20/80

P(reimbursement for a loss is less then 100 and greater then 0)=P(20<X<140)=e-20/80-e-140/80

therefore P(reimbursement for a loss is less then 100 given greater then 0)=(e-20/80-e-140/80)/e-20/80

=0.777

option D